A galvanometer of resistance 20 W gives full scale detlection with a current of 0.004 A. To convert it into an ammeter of range 1 A, the required shunt resistance should be


        (d)    0.38 ohm


        (b)    0.21 ohm


        (c)    0.08 ohm      


(d)    0.05 ohm


Grade:upto college level

1 Answers

Kevin Nash
10 years ago

(c)

Maximum current which can pass through galvanometer,

Ig = 0.004 A

 

Let R be the resistance of shunt.

We know potential drop across AB = Potential drop across CD

R (I -Ig) = Ig (20)

Þ R(1- 0.004) = 0.004 × 20 Þ R = 0.08 W